OpenStudy (anonymous):

I'm trying to use integrals to find the arc length of y=2x-5, -1

OpenStudy (anonymous):

I know the formula and IM getting it too \[\int\limits_{-1}^{3} \sqrt(5) dx\]

OpenStudy (anonymous):

and then i dont know where to go

OpenStudy (anonymous):

\[\sqrt{5} \int\limits_{-1}^{3}dx\]

OpenStudy (anonymous):

just a quick push in the right direction would work great

OpenStudy (anonymous):

ok so what do i do with that?

OpenStudy (anonymous):

is that like taking the integral of 1

OpenStudy (anonymous):

yeah pretty much.

OpenStudy (anonymous):

ok find the derivative dy/dx=2 arc length is

OpenStudy (anonymous):

or you could just jump in and give him the entire thing without helping him through it @sami-21

OpenStudy (anonymous):

k im about to post a really tricky one of these up if someone wont mind helping me again

OpenStudy (anonymous):

i think it is pretty helpful .and i am the first who started it i was typing.i did not jump in.